OpenStudy (anonymous):

Find dy/dx if y=int(t^2+5t-3t)dt interval [2x, 0]

OpenStudy (anonymous):

\[\int\limits_{0}^{2x}t^{2}+5t-3\]

OpenStudy (kc_kennylau):

Okay, so let's deal with the integral first

OpenStudy (anonymous):

that would be t^3/3+5t^2/2-3t

OpenStudy (kc_kennylau):

Don't forget the limits!

OpenStudy (anonymous):

what to do with them?

OpenStudy (kc_kennylau):

\[\Large\left[\frac{t^3}3+5\frac{t^2}2-3t\right]_0^{2x}\]

OpenStudy (kc_kennylau):

Now let's do the substitution

OpenStudy (anonymous):

[(2x)^3/3+5(2x)^2/2-3(2x)]-0

OpenStudy (kc_kennylau):

Expand it

OpenStudy (anonymous):

how lol

OpenStudy (kc_kennylau):

(2x)^3 = 8x^3

OpenStudy (anonymous):

oh like simplify?

OpenStudy (kc_kennylau):

Yep sorry

OpenStudy (tkhunny):

What? Let's not do that. Generally speaking, we have: If \(\dfrac{d}{dt}F(t) = f(t)\) Then, \(\int\limits_{a}^{b}f(t)\;dt = F(b) - F(a)\) Then, \(\int\limits_{0}^{2x}f(t)\;dt = F(2x) - F(0)\) Then, \(\dfrac{d}{dx}\int\limits_{0}^{2x}f(t)\;dt = \dfrac{d}{dx}\left(F(2x) - F(0)\right) = f(2x)\cdot 2 = 2\left((2x)^{2}+5(2x)−3\right)\) The significance of this is that we need only THEORIZE that \(F(t)\) exists. We don't actually have to find \(F(t)\). In addition, there's a lot less to go wrong.